OpenStudy (anonymous):

find derivative: y=(9x+4)^x

rvc (rvc):

can u tell me in which form is it? @hardequalsmath

OpenStudy (anonymous):

what do you mean? i know that you start by taking ln to both sides

rvc (rvc):

listen do u know composite derivates

OpenStudy (anonymous):

i dont think so...

rvc (rvc):

it is in the form a^x?

rvc (rvc):

@hardequalsmath

OpenStudy (anonymous):

no its not the definition formula

rvc (rvc):

y=a^xloga

rvc (rvc):

x^x?

OpenStudy (anonymous):

ok so i know that you start by doing lny=x*ln(9x+4) then take derivative

rvc (rvc):

no

OpenStudy (praxer):

follow the third formula and than again derivative (9x+4)

rvc (rvc):

praxer is correct @hardequalsmath

OpenStudy (anonymous):

isnt that what i did?

rvc (rvc):

yes but u did not mention the derivate of (9x+7)

OpenStudy (anonymous):

cause i didnt take the derivative yet...i was just showing how to set it up

rvc (rvc):

ok sorry

OpenStudy (anonymous):

so then i got y1/y=ln(9x+4)+9x/(9x+4) then what?

rvc (rvc):

@praxer will u plz... check?

OpenStudy (praxer):

i suppose on the left side it should be y'/y cause y*1/y shall yield to 1..Right ?

OpenStudy (anonymous):

yes

OpenStudy (praxer):

Than it is correct. but since you are asked with the y' so keep the y' value on the left and get the final equation. :)

rvc (rvc):

thanks